OpenStudy (anonymous):

find the equation of the line trough the given points.(3, 1) and (1, 3) state the answer in y=mx+b

OpenStudy (anonymous):

y = -m + c

OpenStudy (stormfire1):

Slope m = (y-y1)/(x-x1) so: (3-1)/(1-3) = -1 y = (-1)x+4

OpenStudy (anonymous):

y = -x + 4

OpenStudy (anonymous):

can you do all the steps so I can understand what Im doing wrong

OpenStudy (anonymous):

these are points on a slope

OpenStudy (anonymous):

y-y1=m(x-x1) This one

OpenStudy (anonymous):

1 = 3m + C - equation 1 3 = m + C - equation 2 rearrange eq 2, u will get m = 3 - C substitute this into equation 1 then you will get 1 = 3(3-C) + C 1 = 9 - 3C + C 1 - 9 = -3C + C -8 = -2C C = 4 then substitute C = 4 into general equation of a straight line you will get y = mx + 4 then substitute any point (3,1) or (1,3) into y = mx + 4 i use (3,1) 1 = 3m +4 1-4 = 3m -3 = 3m m = -1 substitute value C and m. y = mx + C y = (-1)x + 4.

OpenStudy (anonymous):

I guess i was totallly wrong. Thanks this makes more scence to me then the way I was trying to do it.

OpenStudy (anonymous):

I don't use the point slope formula or do I use the equation of a line trough two points

OpenStudy (stormfire1):

You already have 2 points so I'd just plug those into the point-slope formula to find m: (y-y1) = m(x-x1) It doesn't matter which pair of coordinates you choose to be x,y or x1,y1...as long as you are consistent throughout the problem. For my example, I'll go with x,y = (1,3) and x1,y1 = (3,1). So it's: (3-1) = m(1-3) To find m, divide both sides by (1-3). You'll end up with m = -1. Now you just need to convert the answer to slope-intercept form which is: y = mx + b So it's: 3 = (-1)1 + b b must = 4

OpenStudy (stormfire1):

Which gives you y = -1x + 4

OpenStudy (anonymous):

ok i totally get that much easier thanks.